when earth and moon are separated by a distance of 3.84x10^8 meters, the magnitude of gravitational force of attraction between them is 2.0x10^20netwons.What would be magnitude of gravitational force if earth and moon separated by distance of 1.92x10^8 meters?
1
Expert's answer
2012-12-19T12:02:26-0500
F=Gm1m2/r^2
r0=3.84x10^8 meters
r1=1.92x10^8 meters
F0=Gm1m2/r0^2
F1=Gm1m2/r1^2
F1=F0(r0^2/r1^2)=F0*4=8.0x10^20netwons